# Auditing Google’s AI Overviews and Featured Snippets: A Case Study on Baby Care and Pregnancy

### v2

@@ Results
In addition to the AIO answers discussed in other Result sections that are visible to the users, we also find 157 AIOs that are generated but not shown to the users. Thus, these AIOs were suppressed by Google. We have excluded these AIOs from our main analyses since they were not visible to the users. However, we have conducted a comparison between the suppressed and the visible AIOs, and established that the suppressed AIOs were characterized by lower relevance than the visible ones and by higher inconsistency. Specifically, while only about 1% of visible AIOs were of low relevance to the query, this was the case for 16% of suppressed ones. In addition, only 31% of suppressed whole AIO answers (versus 67% visible whole AIO answers) were consistent with the FS when both were included in the SERP (see Figure 3 for details). Qualitative analyses of the suppressed AIOs showed that these elements often included either pro-life/anti-abortion or anti-vaccination content.
@@
96.6% of AIOs and 88.7% of FS are rated as highly relevant to the search queries. Only about 1% of AIO/FS showed low relevance to the queries. The detailed percentages for high/medium/low relevance are reported in Table 2. Similarly, in the absolute majority of cases (95.5% for AIO, 89.5% for FS) the highlighted part of text in the AIO/FS is the one most relevant to the query. In Table 4 in the Appendix, we present examples of low relevance ratings for both the full answer and the highlighted text.
@@
For AIO, Chi-square tests indicate a statistically significant difference in the presence of safeguard cues across different question types ($p<0.001$), but not across query sentiments ($p=0.095$). For FS, no statistically significant differences is found in the presence of safeguard cues across either question type or sentiment ($p=0.667$ and $p=0.393$, respectively). While there’s no significant difference among different question sentiments, the fraction of safeguard cue within the set of queries formulated with different sentiment valence is higher than the set of queries formulated as different question types overall , eg., among the set of queries reformulated with sentiment, the fraction range of explicit safeguard cue is 13% to 21% for AIO and 4% to 10% for FS, as shown in Figure 5 (b); conversely, the fraction range of explicit safeguard cue is 0 to 9% for AIO and 0 to 2% for FS among the set of queries with different question types as shown in Figure 5 (a)
@@
In Figure 6 we show the credibility distribution of sources among the top 10% most frequently sourced domains (N=144) in either AIO/FS/10 blue links. Overall, sources with low credibility in the context of health and pregnancy, such as social media and e-commerce websites, are featured comparatively often, ranging from 9% for FS to 17% in AIO and 35% for ten blue links. Low- and medium-credibility sources combined account for nearly half of all citations in AIO and FS. This raises concerns about the reliability of health information presented to users.
@@
RQ4: Query and Response Sentiment Valence
@@
Discussion
Over the past decade, Google’s SERP has evolved from ten blue links to sophisticated search features including top stories, knowledge panels, and, more recently, FS, and AIO. Correspondingly, user behavior has shifted from traditional search patterns (i.e., clicking through multiple vertical links) toward engaging with direct answers that address search intentions immediately, reducing search effort and improving user satisfaction (Wu et al. 2020; Chilton and Teevan 2011). This evolution reflects a consistent trend: aggregating results from heterogeneous sources within a single interface to directly address searcher intent (Oliveira and Teixeira Lopes 2023). While specific SERP features may change, future developments will likely include such as conversational functions (already adopted by other search engines (Capra and Arguello 2023)), and all these will most likely follow this same trajectory of providing immediate answers. As new search features emerge, our research methodology for examining how query characteristics affect SERP results through six quality evaluation metrics will therefore remain relevant and transferable to other high-stakes domains.
Within this evolving landscape, our analysis of baby care and pregnancy searches reveals several concerning implications. The high prevalence of AIO at the top of the SERP, highlights their growing role in shaping how people access and consume information. As AIOs increasingly mediate knowledge access, their quality becomes crucial across other domains like legal advice and political information. Importantly, both query sentiment and question type are significantly associated with the presence of AIOs and FSs, suggesting that seemingly subtle variations in how users formulate their questions may shape which information components they see, especially for featured snippets. This has direct implications for user equity and information exposure across varied demographic or informational contexts, where diverse groups may formulate the same questions differently regarding sentiment and question types.
@@ Discussion
Our study has three main limitations. First, data collection is restricted to the US and English-language results. Second, our analysis represents a single point in time rather than tracking changes longitudinally. Third, we focused exclusively on Google – a choice explained by its dominant market position both in the US and globally over the past decade (Joseph Johnson ; StatCounter ).
@@ Acknowledgments
The work of Aleksandra Urman, Elsa Lichtenegger, and Aniko Hannak was supported by Swiss National Science Foundation Grant 215354; and the work of Robin Forsberg was supported by the Kone Foundation.
@@ Appendix A Geographical location robustness
To investigate whether the geographic location of the user might affect the appearance and content of Featured Snippets (FS) and AI Overviews (AIO), we collected SERPs from five different US cities (New York, NY; Kansas City, MO; San Francisco, CA; Houston, TX; Madison, WI). This additional data collection was conducted during September 2-9, 2024, using a dataset of 528 baby care and pregnancy queries and their variations. The queries were obtained from (1) a researcher who was pregnant and became a new mother, and (2) observational trace data collected from a panel of US residents that includes their web browsing into Google Search. [^6] While this data collection was approved by the institutional IRB, we cannot make the search query dataset from this robustness experiment publicly available because the nature of the data precludes deidentification; therefore, only aggregated results are presented in this manuscript. Since proxy servers and VPNs may not accurately reflect location in Google’s search engine result pages (SERPs) during our manual examination, we used a geospoofing method by injecting precise latitude and longitude coordinates into automated scripts, similar to methods used in previous work (Jung, Juneja, and Mitra 2025).

### v3

@@ 4.1 RQ1: Prevalence of AIO and FS
In addition to the AIO answers discussed in other Result sections that are visible to the users, we also find 157 AIOs that are generated but not shown to the users. Thus, these AIOs were suppressed by Google. We have excluded these AIOs from our main analyses since they were not visible. However, we have conducted a comparison between the suppressed and the visible AIOs, and established that the suppressed AIOs were characterized by lower relevance than the visible ones and by higher inconsistency. Specifically, while only about 1% of visible AIOs were of low relevance to the query, this was the case for 16% of suppressed ones. In addition, only 31% of suppressed whole AIO answers (versus 67% visible whole AIO answers) were consistent with the FS when both were included in the SERP (see Figure 3 for details). Qualitative analyses of the suppressed AIOs showed that these elements often included either pro-life/anti-abortion or anti-vaccination content.
@@ 4.3 RQ3a: Output Relevance
96.6% of AIOs and 88.7% of FS are rated as highly relevant to the queries. Only about 1% of AIO/FS showed low relevance to the queries. The detailed percentages for high/medium/low relevance are reported in Table 2. Similarly, in the absolute majority of cases (95.5% for AIO, 89.5% for FS) the highlighted part of text in the AIO/FS is the one most relevant to the query. In Table 4 in the Appendix, we present examples of low relevance ratings for both the full answer and the highlighted text.
@@ 4.4 RQ3b: Presence of Safeguards
For AIO, Chi-square tests indicate a statistically significant difference in the presence of safeguard cues across different question types ($p<0.001$), but not across query sentiments ($p=0.095$). For FS, no statistically significant differences is found in the presence of safeguard cues across either question type or sentiment ($p=0.667$ and $p=0.393$, respectively). While there’s no significant difference among different question sentiments, the fraction of safeguard cue within the set of queries formulated with different sentiment valence is higher than the set of queries formulated as different question types overall , eg., among the set of queries reformulated with sentiment, the fraction range of explicit safeguard cues is 13% to 21% for AIO and 4% to 10% for FS, as shown in Figure 9 (b) in the Appendix; conversely, the fraction range of explicit safeguard cue is 0 to 9% for AIO and 0 to 2% for FS among the set of queries with different question types as shown in Figure 9 (a).
@@ 4.5 RQ3c: Source Categories
In Figure 5 we show the credibility distribution of sources among the top 10% most frequently sourced domains (N=144) in either AIO/FS/ten blue links. Overall, sources with low credibility in the context of health and pregnancy, such as social media and e-commerce websites, are featured comparatively often, ranging from 9% for FS to 17% in AIO and 35% for ten blue links. Low- and medium-credibility sources combined account for nearly half of all citations in AIO and FS. This raises concerns about the reliability of health information presented to users.
@@
4.6 RQ4: Query and Response Sentiment Valence
@@
5 Discussion
Over the past decade, Google’s SERP has evolved from ten blue links to sophisticated search features including top stories, knowledge panels, and, more recently, FS, and AIO. Correspondingly, user behavior has shifted from traditional search patterns (i.e., clicking through multiple vertical links) toward engaging with direct answers that address search intentions immediately, reducing search effort and improving user satisfaction (Wu et al. 2020). This evolution reflects a consistent trend: aggregating results from heterogeneous sources within a single interface to directly address searcher intent (Oliveira and Teixeira Lopes 2023). While specific SERP features may change, future developments will likely include such as conversational functions (already adopted by other search engines (Capra and Arguello 2023)), and all these will most likely follow this same trajectory of providing immediate answers. As new search features emerge, our research methodology for examining how query characteristics affect SERP results through six quality evaluation metrics will therefore remain relevant and transferable to other high-stakes domains.
Within this evolving landscape, our analysis of baby care and pregnancy searches reveals several concerning implications. The high prevalence of AIO at the top of the SERP highlights their growing role in shaping how people access and consume information. As AIOs increasingly mediate knowledge access, their quality becomes crucial across other domains like legal advice and political information. Importantly, both query sentiment and question type are significantly associated with the presence of AIOs and FSs, suggesting that seemingly subtle variations in how users formulate their questions may shape which information components they see, especially for featured snippets. This has direct implications for user equity and information exposure across varied demographic or informational contexts, where diverse groups may formulate the same questions differently regarding sentiment and question types.
@@ 5 Discussion
Our study has three main limitations. First, data collection is restricted to the US and English-language results. Second, our analysis represents a single point in time rather than tracking changes longitudinally. Third, we focus exclusively on Google – a choice explained by its dominant market position both in the US and globally over the past decade (Joseph Johnson 2021, 2022).
@@ Acknowledgments
The work of Aleksandra Urman, Elsa Lichtenegger, and Aniko Hannak was supported by Swiss National Science Foundation (SNSF) Grant 215354; Joachim Baumann is supported by SNSF grant P500-2_235328; and the work of Robin Forsberg was supported by the Kone Foundation.
@@ Appendix A Geographical location robustness
To investigate whether the geographic location of the user might affect the appearance and content of Featured Snippets (FS) and AI Overviews (AIO), we collected SERPs from five different US cities (New York, NY; Kansas City, MO; San Francisco, CA; Houston, TX; Madison, WI). This additional data collection was conducted during September 2-9, 2024, using a dataset of 528 baby care and pregnancy queries and their variations. The queries were obtained from (1) a researcher who was pregnant and became a new mother, and (2) observational trace data collected from a panel of US residents that includes their web browsing into Google Search. [^6] While this data collection was approved by the institutional IRB, we cannot make the search query dataset from this robustness experiment publicly available because the nature of the data precludes deidentification; therefore, only aggregated results are presented in this manuscript. Since proxy servers and VPNs may not accurately reflect location in Google’s search engine result pages (SERPs) during our manual examination, we used a geospoofing method by injecting precise latitude and longitude coordinates into automated scripts, similar to methods used in previous work (Jung et al. 2025).